Fermi National Laboratory
 

Operation Summary for May 3, 2010 - May 5, 2010

 

Monday May 3

The day shift began with Operations monitoring the Tevatron (TeV) store 7784 (91.8E30) and the Recycler stash (286.1E10), with the Antiproton Source (Pbar) stacking (17.1E10), with the Linac, Booster, and Main Injector (MI) operating normally, with NuMI off due to extraction kicker power supply problems, with the MiniBooNE and Meson MTest experiments taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) experiments using the colliding beams.

During the day shift, Operations terminated store 7784 at 10 AM. D0 and CDF accessed their collision halls and were out by noon. A TeV quench at sector A4 and the B0 low beta magnet occurred during the turn-around. The cryo system recovery was underway by 2:45 PM. Operations put the TeV through a dry squeeze. NuMI extraction kicker repairs continued from the previous shift. NuMI personnel accessed their target hall and the absorber enclosures. The high voltage power supply for an Accumulator kicker (A:IKIK) failed.

During the evening shift, Pbar personnel repaired the A:IKIK power supply and beam returned to Pbar at 4:50 PM. TeV personnel reentered angle bumps for B0 and D0. Operations put the TeV into HEP shot setup at 5:45 PM and established store 7789 at 8:10 PM with an average initial luminosity (AIL) of 270.0E30. NuMI personnel completed their kicker power supply repairs at 8:45 PM.

Tuesday May 4

The midnight shift began with Operations monitoring store 7789 (134.55E30) and the stash (155.8E10), with Pbar stacking (15.8E10), and with NuMI and MiniBooNE taking beam.

During the midnight shift, operators adjusted the NuMI kicker voltage in order to increase the kicker magnet current.

During the day shift, at 9:10 AM, store 7789 aborted and then quenched when the sector D door interlocks went unsecured due to a bad door switch at D2. Once repairs were completed, TeV went through the normal turn-on procedure. MI personnel checked their coalescing. Operations put the TeV into shot setup at 2:45 PM and established store 7792 at 3:50 PM with an AIL of 358.0E30.

During the evening shift, Operations reestablished beam to MTest at 4:45 PM. Operators searched and secured MI-31 early on the evening shift and had the Pelletron operating shortly thereafter. A Booster RF station (BRF16) had a bias supply fault; operators bypassed the station and brought up BRF6 in its place. The NuMI beam permit tripped off at 6 PM. Operations disabled the MTest beam at 6:10 PM. Operators traced the cause of the NuMI trip to a helium leak. Operations contacted NuMI personnel who came in and accessed MI-65 to search for the leak. NuMI personnel reduced the helium flow and resumed sending beam to NuMI at 10:40 PM.

Wednesday May 5

The midnight shift began with Operations monitoring store 7792 (111.08E30) and the stash (293.1E10), with Pbar stacking (11.0E10), and with NuMI and MiniBooNE taking beam.

During the midnight shift, Operations established beam to MTest at 4:25 AM and then disabled it at 5:05 AM for the shot setup. Operations terminated store 7792 at 5:20 AM and then put the TeV into shot setup. Operations established store 7793 at 6:45 AM with an AIL of 397.21E30. Operations reestablished beam to MTest.

The Plans for Wednesday

The plans for today are stack, stash, and store. The Run Coordinator will allow Preaccelerator personnel to access the I- Source dome, for Booster personnel to repair BRF16, and for NuMI personnel to make an access.

 

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 05/05/2010    email Fermilab
Security, Privacy, LegalFermi National Accelerator Laboratory